1: Calculate the wavelength of the fifth line in the hydrogen atom Paschen series?

2: Calculate the energy released in the transition described in the question above?

3: For the hydrogen atom, calculate the wavelength of the light needed to excite the electron from the fifth energy level to the ninth energy level.
Answer #1

DE = (911x10 31 ) x (1.6028101934 8 +(885*10-12 ): (6626410-24; 2* GG282) 2.18 810 18 x 55 J lor, 13.6* 51 16 eu 576 = 2.08*10.17 DE 1 The wavelength : of the line in the spectrum of atomic hydrogen may be described by the Simple relation. * = RH(2 h